Given forecast errors of 4, 8, and -3, what is the mean absolute deviation?

A)4

B)3

C)5

D)6

E)12
Answer #1

1. C. 5

FORECAST ERRORS = 4, 8, -3

SUM(ABSOLUTE ERRORS) = 4 + 8 + 3 = 15
N = 3

MAD = SUM(ABS ERROR) / N = 15 / 3 = 5